# QuantVSR: Low-Bit Post-Training Quantization for Real-World Video Super-Resolution

### v1

@@ Abstract
Diffusion models have shown superior performance in real-world video super-resolution (VSR). However, the slow processing speeds and heavy resource consumption of diffusion models hinder their practical application and deployment. Quantization offers a potential solution for compressing the VSR model. Nevertheless, quantizing VSR models is challenging due to their temporal characteristics and high fidelity requirements. To address these issues, we propose QuantVSR, a low-bit quantization model for real-world VSR. We propose a spatio-temporal complexity aware (STCA) mechanism, where we first utilize the calibration dataset to measure both spatial and temporal complexities for each layer. Based on these statistics, we allocate layer-specific ranks to the low-rank full-precision (FP) auxiliary branch. Subsequently, we jointly refine the FP and low-bit branches to achieve simultaneous optimization. In addition, we propose a learnable bias alignment (LBA) module to reduce the biased quantization errors. Extensive experiments on synthetic and real-world datasets demonstrate that our method obtains comparable performance with the FP model and significantly outperforms recent leading low-bit quantization methods. Code is available at: https://github.com/bowenchai/QuantVSR.
@@ Introduction
Video super-resolution (VSR) is a crucial task in computer vision, aiming to reconstruct high-resolution (HR) videos from low-resolution (LR) inputs. Early VSR methods (Jo et al. 2018; Wang et al. 2019; Nah et al. 2019; Chan et al. 2021; Liang et al. 2024) primarily deal with simple synthetic degradations (e.g., bicubic). However, models trained on such degraded data perform poorly in restoring practical video. In real-world scenarios, videos often suffer from various complex and unknown degradations, which significantly increase the difficulty of video restoration. Numerous methods have been proposed (Yang et al. 2021; Pan et al. 2021; Xie et al. 2023) to address this challenge. Among them, generative adversarial networks (GANs) have gathered attention for their ability to restore fine details (Goodfellow et al. 2014; Lucas et al. 2019). Nevertheless, they still face challenges such as over-smoothing and training instability.
@@
Figure 2: Performance comparisons on the real-world benchmark (i.e., MVSR4x (Wang et al. 2023)) and compression ratio of our method. Our method QuantVSR substantially compresses the model but retains performance comparable to the FP model, which surpasses existing quantization methods (e.g., ViDiT-Q (Zhao et al. 2024) and SVDQuant (Li et al. 2025a)).
@@
To tackle these issues, quantization stands out as an effective compression method (Nagel et al. 2021; Liu et al. 2025). Model quantization maps the floating-point weights and/or activations to low-precision representations, typically as integers, thus decreasing the computational and memory costs. Prior studies have demonstrated the impressive results of model quantization in image/video generation (Li et al. 2023a; Zhao et al. 2024; Li et al. 2025a) and image restoration (Zhu et al. 2025) tasks. However, in VSR, model quantization has rarely been explored, especially for the diffusion-based models. There are two critical difficulties in low-bit quantization for VSR models: (1) Temporal consistency loss. Model quantization introduces inconsistent errors across frames, leading to a loss of temporal coherence in the final generated video. This inconsistency disrupts the smooth transition between consecutive frames, degrading the overall video quality. (2) Complex data distribution. VSR models embed temporal dynamics into latent space representations, resulting in more intricate activation distributions. This complexity increases the challenge of quantization, as both spatial and temporal features must be accounted for to narrow the performance gap between full-precision (FP) and its quantized counterpart.
@@
Quantization is an effective method for model compression, as evidenced by its widespread and successful applications in the field of large language models (Hu et al. 2022; Liu et al. 2023; Xiao et al. 2023; Shao et al. 2023). With the rapid evolution of diffusion models, researchers have concentrated on enhancing their efficiency by applying quantization techniques. PTQ4DM (Shang et al. 2023) pioneers the study of quantized diffusion models, achieving 8-bit quantization. Subsequent works have made some breakthroughs with methods such as specialized calibration strategies (Li et al. 2023a; He et al. 2023b), sensitivity analysis (Yang et al. 2023), quantization-aware training (Li et al. 2023b), weight-only quantization (Sui et al. 2024), and outlier removal (Zhao et al. 2024). Recently, low-bit quantization has made significant progress. EfficientDM (He et al. 2023a) proposes a low-rank quantization fine-tuning strategy. While SVDQuant (Li et al. 2025a) employs a 16-bit parallel low-rank branch. Both EfficientDM and SVDQuant achieve 4-bit quantization with minimal performance loss. Furthermore, there are methods designed specifically for particular tasks and models. such as PassionSR (Zhu et al. 2025) for one-step diffusion-based super-resolution models, ViDiT-Q (Zhao et al. 2024) for image and video generation. However, studies on the quantization of VSR models remain limited. Existing quantization methods cause significant performance loss when applied to VSR models.
@@
The quantization calibration process consists of three main stages. First, we analyze the spatio-temporal complexity distribution of the calibration dataset. Such calibration statistics allocate higher ranks to the FP branches in layers with high complexity. Second, we refine two low-rank matrices and residual, allowing the STCA structure to attain optimal performance across both FP and low-bit branches in each layer. Finally, as changes in quantized weights introduced by previous modules affect the biased quantization errors, the LBA module is trained in the last stage with all other parameters frozen. For fine-tuning and training in the final two stages, we adopt the mean squared error (MSE) between FP and quantized model outputs as the loss function.

### v2

@@ Abstract
Diffusion models have shown superior performance in real-world video super-resolution (VSR). However, the slow processing speeds and heavy resource consumption of diffusion models hinder their practical application and deployment. Quantization offers a potential solution for compressing the VSR model. Nevertheless, quantizing VSR models is challenging due to their temporal characteristics and high fidelity requirements. To address these issues, we propose QuantVSR, a low-bit quantization model for real-world VSR. We propose a spatio-temporal complexity aware (STCA) mechanism, where we first utilize the calibration dataset to measure both spatial and temporal complexities for each layer. Based on these statistics, we allocate layer-specific ranks to the low-rank full-precision (FP) auxiliary branch. Subsequently, we jointly refine the FP and low-bit branches to achieve simultaneous optimization. In addition, we propose a learnable bias alignment (LBA) module to reduce the biased quantization errors. Extensive experiments on synthetic and real-world datasets demonstrate that our method obtains comparable performance with the FP model and significantly outperforms recent leading low-bit quantization methods.
@@ Introduction
Video super-resolution (VSR) is a crucial task that aims to reconstruct high-resolution (HR) videos from low-resolution (LR) inputs. Early VSR methods (Jo et al. 2018; Wang et al. 2019; Nah et al. 2019; Chan et al. 2021; Liang et al. 2024) assume synthetic degradations (e.g., bicubic). However, models trained on such degraded data perform poorly in restoring practical videos. In real-world scenarios, videos often suffer from various complex and unknown degradations, which increase the difficulty of video restoration. Numerous methods are proposed (Yang et al. 2021; Pan et al. 2021; Xie et al. 2023) to address this challenge. Among them, generative adversarial networks (GANs) gather attention for their ability to restore fine details (Goodfellow et al. 2014; Lucas et al. 2019). Nevertheless, they still face challenges such as over-smoothing and training instability.
@@
Figure 2: Performance comparisons on the real-world benchmark (i.e., MVSR4x (Wang et al. 2023)) and compression ratio of our method. Bn represents n-bit quantization. Our QuantVSR compresses the model but retains performance comparable to the FP model, which surpasses existing quantization methods (e.g., ViDiT-Q (Zhao et al. 2024) and SVDQuant (Li et al. 2025a)).
@@
To tackle these issues, quantization stands out as an effective compression method (Nagel et al. 2021; Liu et al. 2025). Model quantization maps floating-point weights and/or activations to low-precision representations, typically as integers, thereby decreasing computational and memory costs. Prior studies have demonstrated the impressive results of quantization in image/video generation (Li et al. 2023a; Zhao et al. 2024; Li et al. 2025a) and image restoration (Zhu et al. 2025) tasks. However, in VSR, model quantization has rarely been explored, especially for the diffusion-based models. There are two critical difficulties in low-bit quantization for VSR models: (1) Temporal consistency loss. Model quantization introduces inconsistent errors across frames, resulting in a loss of temporal coherence in the final generated video. (2) Complex data distribution. VSR models embed temporal dynamics into latent features, resulting in more intricate activation distributions. This increases quantization challenges, as both spatial and temporal features must be considered to narrow the performance gap between full-precision (FP) and its quantized counterpart.
@@
Quantization is an effective method for model compression, as evidenced by its widespread and successful applications in the field of large language models (Hu et al. 2022; Liu et al. 2023; Xiao et al. 2023; Shao et al. 2023; Huang et al. 2024). With the rapid evolution of diffusion models, researchers have concentrated on enhancing their efficiency by applying quantization techniques. PTQ4DM (Shang et al. 2023) pioneers the study of quantized diffusion models, achieving 8-bit quantization. Subsequent works have made breakthroughs with methods such as specialized calibration strategies (Li et al. 2023a; He et al. 2023b), sensitivity analysis (Yang et al. 2023), quantization-aware training (Li et al. 2023b), weight-only quantization (Sui et al. 2024), and outlier removal (Zhao et al. 2024). Recently, low-bit quantization has made significant progress. EfficientDM (He et al. 2023a) proposes a low-rank quantization fine-tuning strategy. While SVDQuant (Li et al. 2025a) employs a 16-bit parallel low-rank branch. Both EfficientDM and SVDQuant achieve 4-bit quantization with minimal performance loss. Furthermore, there are methods designed specifically for particular tasks and models. such as PassionSR (Zhu et al. 2025) for one-step diffusion-based super-resolution models, ViDiT-Q (Zhao et al. 2024) for image and video generation. However, studies on the quantization of VSR models remain limited. Existing quantization methods cause significant performance loss when applied to VSR models.
@@
The quantization calibration process consists of three main stages. First, we analyze the spatio-temporal complexity distribution of the calibration dataset. Such calibration statistics allocate higher ranks to the FP branches in layers with high complexity. Second, we refine two low-rank matrices and residual, allowing the STCA structure to attain optimal performance across both FP and low-bit branches in each layer. Finally, as changes in quantized weights introduced by previous modules affect the biased quantization errors, the LBA module is trained in the last stage with all other parameters frozen. For fine-tuning and training in the final two stages, we adopt the mean squared error (MSE) between FP and quantized model outputs (i.e., $\boldsymbol{z}_{fp}$ and $\boldsymbol{z}_{q}$).
@@ Conclusion
Acknowledgments
This work was supported by Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and the Fundamental Research Funds for the Central Universities.